Line PR is perpendicular to line QS, angle QPR=60°, angle PSR=30° and line PQ=8cm. Calculate line SR correct to two significant figures
1
Expert's answer
2020-06-28T18:02:36-0400









The information given could well understood if put into a sketch as shown above thus we get to know that PQ = PS = 8, QPR = RPS=60, PSR=PQR = 30

TO calculate SR we use sine rule, ps/sin SRP= SR/sin SPR THUS,

8/Sin90= SR/sin60, SR= (sin60*8)/sin90 , SR = 6.93 , SR = 6.9